OpenStudy (jenniferjuice):

Solve for x: |3x + 2| = 2

OpenStudy (jenniferjuice):

x = 0 and x = - 4 / 3 x = 4 / 3 and x = - 4 / 3 x = 0 No Solutions

OpenStudy (jenniferjuice):

@blue_cheetah @Bearsfan98 @bigmatt can you guys help me out please?

OpenStudy (goldphenoix):

What's te absolute value of 3x+2 ?

OpenStudy (jenniferjuice):

i have no idea

OpenStudy (goldphenoix):

So you have 0% knowledge of Absolute Value? ex: |-3| You know that absolute value of -3?

OpenStudy (jenniferjuice):

oh yeah nvm i remember

OpenStudy (goldphenoix):

So what's the absolute value of 3x + 2?

OpenStudy (jenniferjuice):

2? maybe

OpenStudy (goldphenoix):

Well. The absolute value of 2 is 2. But what's the absolute value of 3x?

OpenStudy (anonymous):

I would think it's 3|x|. The 3 can come out, but the x can be anything still. That is if you are only looking at the |3x|

OpenStudy (jenniferjuice):

3? @GoldPhenoix

OpenStudy (goldphenoix):

Nope. I'll explain this to you later. The absolute value of 3x + 2 is 3x + 2. So: |3x + 2| = 3x+2 So now your equation looks like: 3x + 2 = 2 Now solve it as an ordinary equation. :)

OpenStudy (jenniferjuice):

so 3x=0

OpenStudy (goldphenoix):

Hmm. Wait.

OpenStudy (jenniferjuice):

pleaze hurry i have to finish by 12;00 and i have 9 min !!!

OpenStudy (anonymous):

1. First eliminate the absolute value (by rewriting this is the rule) \[3x+2=2\]\[3x = 0\]\[x = 0\] Now you need to do the equation that is made by rewriting. \[3x+2= -2\]\[3x=-4\]\[x=-\frac{ 4 }{ 3 }\]

OpenStudy (goldphenoix):

^Right on the spot.

OpenStudy (anonymous):

So when you are eliminating the absolute value, you should always start by writing 2 equations. In this case we write: \[3x+2=2\]and \[3x+2=-2\] Then solve both.

OpenStudy (jenniferjuice):

x = 0 and x = - 4 / 3 so that?

OpenStudy (jenniferjuice):

?????????????????????????????????????????????????????/

OpenStudy (goldphenoix):

Yes.

OpenStudy (jenniferjuice):

okay thankyou!!!